OpenStudy (anonymous):

Which of the following is a monomial? a.c^2+c+6 b.c^2-16 c.-8c d.c^3+4c^2-12c+7

OpenStudy (z4k4r1y4):

'mono' means one

pooja195 (pooja195):

Which one has ONE term?

OpenStudy (anonymous):

im not to sure how mono works everytime i start or try it i get completely lost!

pooja195 (pooja195):

Mononomial = one term example 8x^2 binomial = 2 terms 78+45x trinomial=3 terms x^2+4v+v polynomial= MANY terms more than 5 6+7x+8x+9x+6x^2+78

pooja195 (pooja195):

does that help?

OpenStudy (anonymous):

yes thanks!!!

pooja195 (pooja195):

:)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

how can you tell polynomial is ascending order?

pooja195 (pooja195):

it would b greatest to least :-) number wise...